OpenStudy (anonymous):

Optimization problem: "A farmer has 140 ft of fencing with which she will enclose a rectangular region adjacent to a barn 60 ft long. If she insists on using the entire length of the barn as a portion of one side of the rectangle, what is the maximum area that may be enclosed?'

OpenStudy (zzr0ck3r):

ok so your constraint is 2l+2w = 140+60 note we have 140 to use and 60 already there and our function is f(x) = l*w

OpenStudy (zzr0ck3r):

@kissy do you understand?

OpenStudy (anonymous):

Kind of

OpenStudy (anonymous):

x + 60 = side of rectangle that includes the barn. y = opposite side of rectangle. z = the length of one of the two other sides. w = area to be maximized.

OpenStudy (anonymous):

What I don't understand is the w=yz and the y=x+z part